# Two-dimensional electronic spectroscopy in the condensed phase using equivariant transformer accelerated molecular dynamics simulations

### v1

@@ Abstract
Recent work has suggested opportunities to use machine learning (ML) approaches to provide a tractable route to efficiently compute and understand electronic properties ranging from electronic excitation energies in the gas phase to condensed phase linear and 2D electronic spectra. For gas-phase systems these ML approaches have been used to predict orbital energy gapsFuchs et al. (2020); Luo et al. (2023); Liao and Smidt (2023); Cignoni et al. (2024) and excitation energiesRamakrishnan et al. (2015); Pronobis et al. (2018); Ghosh et al. (2019); Mazouin, Schöpfer, and von Lilienfeld (2022); Entwistle et al. (2023); Baker, Pasini, and Hauck (2024) for databases of molecules at their optimized geometries. To obtain gas phase linear spectra that incorporate the shape of the electronic absorption peaks arising from molecular geometries beyond their optimized structure, recent work has trained ML models to predict electronic energy gaps, i.e. the difference in energy between the electronic states, at multiple geometries and combined these with the ensemble approachMukamel (1995); Crespo-Otero and Barbatti (2012); Isborn et al. (2012); Ge et al. (2015); Marenich, Cramer, and Truhlar (2015); Milanese et al. (2017); Zuehlsdorff et al. (2017) to compute linear spectra Ye et al. (2019); Xue, Barbatti, and Dral (2020); Lu et al. (2020); Westermayr, Marquetand, and Marquetand (2020); Zou et al. (2023); Vinod et al. (2023); Petrusevich et al. (2023). ML potential energy surfaces have also been trained to the quality required to perform nonadiabatic dynamics simulations to study electronic relaxation through conical intersections in gas phase systems Chen et al. (2018); Westermayr et al. (2019); Axelrod, Shakhnovich, and Gómez-Bombarelli (2022); Shakiba and Akimov (2024). Machine learned potentials also show considerable promise in condensed phase studies such as in the construction of model Hamiltonians for large networks of coupled chromophores Häse et al. (2016); Farahvash et al. (2020); Cignoni, Cupellini, and Mennucci (2023); Betti et al. (2024), calculating ensemble approach linear electronic absorption spectra by machine learning electronic energy gaps in implicit solvent Bononi et al. (2020); Chen et al. (2022), and using electrostatic machine learning embedding to accurately capture environmental and vibronic effectsZinovjev and Curutchet (2025).
@@
Here we show that using an Equivariant Transformer (EqT) architecture Vaswani et al. (2017); Thölke and de Fabritiis (2021); Pelaez et al. (2024), which naturally encodes permutational, rotational, and translational symmetries within the machine-learned model, one can capture both the ground state electronic potential energy and electronic energy gap surfaces to simulate atomistic MD, the excitation energy gaps, and the corresponding linear and 2D electronic absorption spectra of a solvated chromophore using the energies and gradients obtained from only 2546 ground state and 90 excited state electronic structure calculations (SI Sec. IB-ID) compared to $\sim$100,000 of both without the use of ML. We demonstrate this approach by simulating the electronic spectroscopy of Nile blue in ethanol and comparing it to recent 2DES experimentsSon, Mosquera-Vázquez, and Schlau-Cohen (2017). In doing so, we reveal how atomistic MD simulations can be used to unravel the rich information present in 2DES signals and highlight the additional understanding that one can extract about the role of the atomistic environment. The EqT approach offers significant improvements over our previous work as an EqT model trained with less than 100 excited state calculations can achieve a root mean squared error (RMSE) of less than 10 meV compared to 1000s of training calculations achieving RMSEs of 50-80 meV using the atom-centered neural network architecture based on Chebyshev polynomial descriptorsArtrith, Urban, and Ceder (2017) that we previously employedChen et al. (2020). The EqT architecture thus offers an extremely data-efficient route to fully atomistic, condensed-phase predictions of linear and 2DES spectra. Avoiding the tour-de-force ab initio molecular dynamics and excited state electronic structure calculations previously required will enable tighter feedback loops between experiment and simulation. Upon establishing the efficiency of this approach, we then compare our simulations of the linear, 2DES, and pump-probe spectra to recent experiments of Nile blue in ethanol to unravel how solvation both drives electronic energy relaxation and enhances coupling to specific chromophore vibrational motions.
@@
To calculate linear and 2DES spectra, we use the second-order cumulant approach, Mukamel (1985, 1995); Zuehlsdorff et al. (2019) which requires obtaining a one-time correlation function of electronic energy gaps from an MD trajectory. To train an ML surface on which to perform the ground state dynamics, we began with a 50 ps density-functional tight-binding (DFTB) trajectory of Nile blue solvated by 175 ethanol molecules, selected configurations, calculated their energies and gradients with DFT using the revPBE exchange-correlation functional, trained an EqT model, and then iteratively added more configurations until the machine-learned potential was stable and achieving a gradient RMSE error on the test set of 44 meV/Å (SI Sec. IB). To calculate the ground to excited state energy difference, we sampled configurations from the ground-state EqT dynamics, calculated CAM-B3LYP/6-31G* excitation energies and excited state gradients, treating the Nile blue and, on average, 118 ethanol molecules at the same level of theory using TeraChem v1.9 Seritan et al. (2021)(SI Sec. IC), and used these values to train a second EqT model. We used the final ground state EqT model to generate 1 ns of dynamics then used the excitation energy EqT model to predict the energy gaps for every configuration in that trajectory, yielding a trajectory of 500,000 energy gap predictions trained using fewer than 200 excited state electronic structure calculations (SI Sec. ID).
@@
Given the converged spectra obtained from the EqT model, we now compare these spectra with recent experiments. The top panel of Fig. 2 shows the simulations and experimentally obtained linear absorption spectrum of Nile blue in ethanolSon, Mosquera-Vázquez, and Schlau-Cohen (2017). To allow easier comparison of the shape of the spectrum, the simulated spectrum has been shifted such that its absorption maximum is aligned with the experiment. The energy gap model was trained using TDDFT with the CAM-B3LYP exchange-correlation functional and 6-31G basis set, which is known to systematically overestimate electronic excitation gaps of organic chromophores by 100s of meV Shao et al. (2019), and in this case overestimates it by 0.32 eV (the spectral maximum is 1.97 eV experimentally and 2.29 eV from simulation). One could improve the description of the position of the absorption maximum by using transfer learning to higher level wavefunction methodsChen et al. (2023a, b), but since our primary focus here is the mechanisms by which solvent interactions give rise to the spectral shape, we apply a uniform 0.32 eV shift of the linear and 2DES spectra to allow for easier comparison with experiment. With the shift applied, our simulated linear absorption spectrum yields good agreement with the experiment (Fig. 2) with a small underestimate of the width and high-frequency tail of the spectrum. The slightly narrower spectrum may be due to TDDFT with CAM-B3LYP underestimating the effect on the excitation energies upon H-bonding with the solvent, which has recently been demonstrated for the green fluorescent protein chromophore in water where EOM-CCSD was shown to correctly capture this effectChen et al. (2023a), or it may be due to non-Condon fluctuations, which have also recently been shown to produce additional broadeningWiethorn et al. (2023).
To better understand the origins of the shape of the linear spectrum, we can analyze the spectral density from which it was obtained within the second-order cumulant expansion of the energy gap fluctuations between the ground and first excited stateMukamel (1985, 1995); Zuehlsdorff et al. (2019). The spectral density encodes how strongly nuclear motions, including both those of the chromophore and its solvation environment, couple to the electronic excitation between the ground and first excited state, and thus provides information about how these motions modulate electronic energy transfer. The simulated spectral density is shown in the bottom panel of Fig. 2 accompanied by the experimentally obtained surface-enhanced Raman scattering (SERS) spectrum of Nile blue in ethanol Auguié et al. (2012). The SERS spectrum and the simulated spectral density correspond to different observables and so the relative intensities of the peaks should not be directly compared but one generally expects to see similar features since they both are sensitive to motions that couple strongly to the electronic excitation. In particular, both have a sharp prominent peak at $\sim$590 $\mathrm{cm}^{-1}$ that corresponds to a ring breathing mode of the central heteroatom-containing ring of Nile blueLawless and Mathies (1992). The other prominent experimentally observed feature, a 1640 $\mathrm{cm}^{-1}$ vibration primarily involving motion of the charged amine, is red-shifted and lower intensity than the corresponding highest frequency SERS peak. A normal mode analysis (SI Sec. III) of the Nile blue molecule in vacuum using the same electronic structure method as that for training the ground state dynamics shows a set of four modes ranging from 1578 to 1636 $\mathrm{cm}^{-1}$ that all contain contributions from the scissor motion of the -NH+${}_{2}^{+}$ (Fig. 2 top, inset). These modes overlap closely with the peak in the simulated spectral density, indicating that the deviation from experiment is primarily a result of the level of electronic structure theory and not errors in the machine learning or effects from solvation. The difference in intensity reflects the different weighting SERS and spectral densities give to vibrational features with the low intensity in the simulated spectrum indicating that this vibrational mode is weakly coupled to the electronic transition.
@@
2DES spectra present a more strenuous test of simulation than linear absorption spectra owing to the richer details they encode. Fig. 3 compares the 2DES spectra obtained from simulation to that from experiment. The simulated spectra are shifted by the same amount used for the linear spectra and we have multiplied the signal by the experimental pulse spectral profile to account for the window of frequencies observable in the experiment (SI Sec. II). The features and general shape of the 2DES obtained from the experiment and simulation are in good agreement, except for the negative feature observed in the experimental signal to the right of the main peak. Such negative features often arise from excited state absorption (ESA) processes but transient absorption experiments suggest that the negative feature observed here is likely an artifact from the phasing process. Lewis et al. (2010); Son, Mosquera-Vázquez, and Schlau-Cohen (2017); Dostál et al. (2018); Mewes et al. (2021) The simulated signal does not have any negative intensity because our simulation only includes the ground and first excited states, so there are no higher electronic states to absorb into. Negative ESA contributions above the main peak in experimental 2DES spectra shift the peak maxima slightly lower along the diagonal. In contrast, this shift does not occur in the simulated spectra that do not capture ESA features. For an additional comparison, one can integrate 2DES spectra along the excitation axis to obtain the pump-probe spectrumMukamel (1995); Jonas (2003) which as shown in SI Sec. IV yields good agreement between experiment and theory.
@@ Acknowledgments
Data Availability
The data that supports the findings of this study are available within the article and its supplementary material.

### v2

@@ Abstract
Recent work has suggested opportunities to use machine learning (ML) approaches to provide a tractable route to efficiently compute and understand electronic properties ranging from electronic excitation energies in the gas phase to condensed phase linear and 2D electronic spectra. For gas-phase systems these ML approaches have been used to predict orbital energy gapsFuchs et al. (2020); Luo et al. (2023); Liao and Smidt (2023); Cignoni et al. (2024) and excitation energiesRamakrishnan et al. (2015); Pronobis et al. (2018); Ghosh et al. (2019); Mazouin, Schöpfer, and von Lilienfeld (2022); Entwistle et al. (2023); Baker, Pasini, and Hauck (2024) for databases of molecules at their optimized geometries. To obtain gas phase linear spectra that incorporate the shape of the electronic absorption peaks arising from molecular geometries beyond their optimized structure, recent work has trained ML models to predict electronic energy gaps, i.e. the difference in energy between the electronic states, at multiple geometries and combined these with the ensemble approachMukamel (1995); Crespo-Otero and Barbatti (2012); Isborn et al. (2012); Ge et al. (2015); Marenich, Cramer, and Truhlar (2015); Milanese et al. (2017); Zuehlsdorff et al. (2017) to compute linear spectra Ye et al. (2019); Xue, Barbatti, and Dral (2020); Lu et al. (2020); Westermayr and Marquetand (2020); Zou et al. (2023); Vinod et al. (2023); Petrusevich et al. (2023). ML potential energy surfaces have also been trained to the quality required to perform nonadiabatic dynamics simulations to study electronic relaxation through conical intersections in gas phase systems Chen et al. (2018); Westermayr et al. (2019); Axelrod, Shakhnovich, and Gómez-Bombarelli (2022); Shakiba and Akimov (2024). Machine learned potentials also show considerable promise in condensed phase studies such as in the construction of model Hamiltonians for large networks of coupled chromophores Häse et al. (2016); Farahvash et al. (2020); Cignoni, Cupellini, and Mennucci (2023); Betti et al. (2024), calculating ensemble approach linear electronic absorption spectra by machine learning electronic energy gaps in implicit solvent Bononi et al. (2020); Chen et al. (2022), and using electrostatic machine learning embedding to accurately capture environmental and vibronic effectsZinovjev and Curutchet (2025).
@@
Here we show that using an Equivariant Transformer (EqT) architecture Vaswani et al. (2017); Thölke and de Fabritiis (2021); Pelaez et al. (2024), which naturally encodes permutational, rotational, and translational symmetries within the machine-learned model, one can capture both the ground state electronic potential energy and electronic energy gap surfaces to simulate atomistic MD, the excitation energy gaps, and the corresponding linear and 2D electronic absorption spectra of a solvated chromophore using the energies and gradients obtained from only 2546 ground state and 90 excited state electronic structure calculations (SI Sec. IB-ID) compared to $\sim$100,000 of both without the use of ML. We demonstrate this approach by simulating the electronic spectroscopy of Nile blue in ethanol and comparing it to recent 2DES experimentsSon, Mosquera-Vázquez, and Schlau-Cohen (2017). In doing so, we reveal how atomistic MD simulations can be used to unravel the rich information present in 2DES signals and highlight the additional understanding that one can extract about the role of the atomistic environment. The EqT approach offers significant improvements over our previous work as an EqT model trained with less than 100 excited state calculations can achieve a root mean squared error (RMSE) of less than 10 meV compared to 1000’s of training calculations achieving RMSEs of 50-80 meV using the atom-centered neural network architecture based on Chebyshev polynomial descriptorsArtrith, Urban, and Ceder (2017) that we previously employedChen et al. (2020). The EqT architecture thus offers an extremely data-efficient route to fully atomistic, condensed-phase predictions of linear and 2DES spectra. Avoiding the tour-de-force ab initio molecular dynamics and excited state electronic structure calculations previously required will enable tighter feedback loops between experiment and simulation (SI Sec. IE). Upon establishing the efficiency of this approach, we then compare our simulations of the linear, 2DES, and pump-probe spectra to recent experiments of Nile blue in ethanol to unravel how solvation both drives electronic energy relaxation and enhances coupling to specific chromophore vibrational motions.
@@
To calculate linear and 2DES spectra, we use the second-order cumulant approach, Mukamel (1985, 1995); Zuehlsdorff et al. (2019) which requires obtaining a one-time correlation function of electronic energy gaps from an MD trajectory. To train an ML surface on which to perform the ground state dynamics, we began with a 50 ps density-functional tight-binding (DFTB)Porezag et al. (1994); Seifert, Porezag, and Frauenheim (1996) trajectory of Nile blue solvated by 175 ethanol molecules, selected configurations, calculated their energies and gradients with DFT using the revPBE exchange-correlation functional, trained an EqT model, and then iteratively added more configurations until the machine-learned potential was stable and achieving a gradient RMSE error on the test set of 44 meV/Å (SI Sec. IB). To calculate the ground to excited state energy difference, we sampled configurations from the ground-state EqT dynamics, calculated CAM-B3LYP/6-31G* excitation energies and excited state gradients, treating the Nile blue and, on average, 118 ethanol molecules at the same level of theory using TeraChem v1.9 Seritan et al. (2021) (SI Sec. IC), and used these values to train a second EqT model. We used the final ground state EqT model to generate 1 ns of dynamics then used the excitation energy EqT model to predict the energy gaps for every configuration in that trajectory, yielding a trajectory of 500,000 energy gap predictions trained using fewer than 200 excited state electronic structure calculations (SI Sec. ID).
@@
Given the converged spectra obtained from the EqT model, we now compare these spectra with recent experiments. The top panel of Fig. 2 shows the simulations and experimentally obtained linear absorption spectrum of Nile blue in ethanolSon, Mosquera-Vázquez, and Schlau-Cohen (2017). To allow easier comparison of the shape of the spectrum, the simulated spectrum has been shifted such that its absorption maximum is aligned with the experiment. The energy gap model was trained using TDDFT with the CAM-B3LYP exchange-correlation functional and 6-31G basis set, which is known to systematically overestimate electronic excitation gaps of organic chromophoresLaurent and Jacquemin (2013); Shao et al. (2019) by 100s of meV, and in this case overestimates it by 0.32 eV (the spectral maximum is 1.97 eV experimentally and 2.29 eV from simulation). One could improve the description of the position of the absorption maximum by using transfer learning to higher level wavefunction methodsChen et al. (2023a, b, c), but since our primary focus here is the mechanisms by which solvent interactions give rise to the spectral shape, we apply a uniform 0.32 eV shift of the linear and 2DES spectra to allow for easier comparison with experiment. With the shift applied, our simulated linear absorption spectrum yields good agreement with the experiment (Fig. 2) with a small underestimate of the width and high-frequency tail of the spectrum. The slightly narrower spectrum may be due to TDDFT with CAM-B3LYP underestimating the effect on the excitation energies upon H-bonding with the solvent, which has recently been demonstrated for the green fluorescent protein chromophore in water where EOM-CCSD was shown to correctly capture this effectChen et al. (2023a), or it may be due to non-Condon fluctuations, which have also recently been shown to produce additional broadeningWiethorn et al. (2023).
To better understand the origins of the shape of the linear spectrum, we can analyze the spectral density, $J(\omega)$, which can be obtained fromMukamel (1985, 1995); Zuehlsdorff et al. (2019) the time correlation of the energy gap fluctuations $\delta U$ relative to the mean electronic energy gap,
$J(\omega)=\frac{\beta\omega}{2}\int_{-\infty}^{\infty}dt~e^{i\omega t}\langle\delta U(t)\delta U(0)\rangle.$ (1)
[…]
@@
2DES spectra present a more strenuous test of simulation than linear absorption spectra owing to the richer details they encode. Fig. 3 compares the 2DES spectra obtained from simulation to that from experiment. The simulated spectra are shifted by the same amount used for the linear spectra and we have multiplied the signal by the experimental pulse spectral profile to account for the window of frequencies observable in the experiment (SI Sec. II). The features and general shape of the 2DES obtained from the experiment and simulation are in good agreement, except for the negative feature observed in the experimental signal to the right of the main peak. Such negative features can arise from excited state absorption (ESA) processes but transient absorption experiments suggest that the negative feature observed here is likely an artifact from the phasing process. Lewis et al. (2010); Son, Mosquera-Vázquez, and Schlau-Cohen (2017); Dostál et al. (2018); Mewes et al. (2021). The simulated signal does not have any negative intensity because our simulation only includes the ground and first excited states, so there are no higher electronic states to absorb into. While the negative features observed at higher excitation frequencies in the experimental 2DES likely arise from phasing artifacts, the experimental pump-probe spectra (SI Sec. IV) show negative intensity that arises from ESA above $\sim 2000~\mathrm{cm}^{-1}$. These pronounced negative intensity contributions from ESA should appear as negative intensity on the 2DES emission axis since the integral of the excitation axis onto it yields the time-dependent pump-probe spectra.Mukamel (1995); Jonas (2003) However, these contributions lie above 2000 $\mathrm{cm}^{-1}$ and hence they are not observed as they lie outside the range of observable frequencies due to the experiment’s pulse spectral profile. Despite this, ESA still affects the experimental 2DES by shifting the position of the peak maximum to a lower frequency than that of the linear absorption spectrum. This shift in the 2DES maximum is not captured in the simulations that do not include ESA.
@@
Supporting Information
Additional details of the experimental methods, electronic structure calculations, training of machine learning models, molecular dynamics simulations, assessment of ML-based computational cost acceleration, normal modes that contribute to the spectral density of Nile blue in ethanol, comparison between experimental and simulated pump-probe spectra, assessment of the GSB and SE components of the simulated 2DES spectra without the laser profile, reweighting procedure to obtain the Stoke shift, and hydrogen bonding sites of Nile blue (PDF)
Data sets, EqT models, and input files to train the ground state and electronic excitation energy gap EqT models (10.5281/zenodo.15376273)